OpenStudy (sup_open~study):

The density of a particular rock is 1.2073 g/cm(3) a. How many grams of the rock are contained in a 1044.55-cm(3) sample of the rock? b. How many cm(3) will be occupied by 278.816 g of the rock?

OpenStudy (sup_open~study):

i have no idea what to do :(

OpenStudy (ajprincess):

density=mass/volume

OpenStudy (ajprincess):

a) 1.2073g/cm^3=mass in gram/1044.55cm^3

OpenStudy (ajprincess):

b) 1.2073g/cm^3=278.816 g /volume in cm^3

OpenStudy (ajprincess):

Can u find it?

OpenStudy (sup_open~study):

a is 1139692528

OpenStudy (sup_open~study):

??

OpenStudy (ajprincess):

hmm no. mass=1.2073*1044.55 =1261.085g

OpenStudy (sup_open~study):

i'm sorry i have no idea on how to do this :(

OpenStudy (ajprincess):

density=mass/volume If u know the values of density and volume and want to find the value of mass u hav to multiply the values of density and volume. mass=density*volume If u know the values of density and mass and want to find the value of volume u hav to divide the value of mass by the value of density . volume=mass/density Is it clear?

OpenStudy (sup_open~study):

yes thank you

OpenStudy (ajprincess):

yw:)

OpenStudy (sup_open~study):

were is the volume at ???

OpenStudy (ajprincess):

sorry didnt get u?

OpenStudy (sup_open~study):

i mean how would i find the values of density ???

OpenStudy (ajprincess):

If u know the values of mass and volume and want to find the value of density u hav to divide the value of mass by the value of volume density=mass/volume

OpenStudy (ajprincess):

Is it clear?

OpenStudy (sup_open~study):

i don't know the values of mass and value shood there be a chart telling you what they are ???

OpenStudy (sup_open~study):

thank you for your help i will try to go before school tomorrow and ask for some help i have a horrible chem teacher pretty much the whole class is failing her including me xD

OpenStudy (ajprincess):

in the a)part of the question u r asked to find the value of mass. the values of density and the volume are given to u. density=1.2073 g/cm^3 volume=1044.55cm^3 To find the value of mass use mass=volume*density =1044.55cm^3*1.2073 g/cm^3 in the b)part of the question u r asked to find the value of volume. the values of density and the mass are given to u. density=1.2073 g/cm^3 mass=278.816 g To find the value of mass use volume=mass/density =278.816 g/1.2073 g/cm^3